OpenStudy (anonymous):

Charlie put postage worth $1.29 on a package he sent to his sister. He used only 16-cent stamps and 7-cent stamps. How many each of type of stamp did he use? Show work and explain it.

OpenStudy (anonymous):

Any help?

OpenStudy (saifoo.khan):

Make 2 equations.

OpenStudy (saifoo.khan):

Wait! Are you sure the question's complete?

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

And I think yes the question is complete that is what is on my paper.

OpenStudy (saifoo.khan):

@dumbcow @Zarkon

OpenStudy (saifoo.khan):

Suppose number of 16 cents stamps as "x" Suppose number of 7 cents stamps as "y" 16x + 7y = 129 - 1 I can't make up the second question.

OpenStudy (saifoo.khan):

@lgbasallote

OpenStudy (anonymous):

Why did you subtract 1 ?

OpenStudy (saifoo.khan):

No, that's sign for equation 1.

OpenStudy (saifoo.khan):

I still think that the question is incomplete.. They should've given you the total number of stamps.

OpenStudy (anonymous):

You are probably right; but do you know what is the exact answer?

OpenStudy (saifoo.khan):

Yes. You can solve it using a shortcut.

OpenStudy (anonymous):

How?

OpenStudy (saifoo.khan):

As now you have the equation, 16x + 7y = 129 Insert a particular value of x and y and make the answer equal to 129,

OpenStudy (saifoo.khan):

Take x = 5, y = 7

OpenStudy (lgbasallote):

this question isn't complete...

OpenStudy (lgbasallote):

it should state how many stamps there are overall

OpenStudy (anonymous):

I know but that Is what my teacher gave me.

OpenStudy (saifoo.khan):

The stamps are: 5+7 = 12

OpenStudy (anonymous):

But you can see figure it out because at least the gave you the cents for each stamp and the total stamps.

OpenStudy (anonymous):

So there is 12 stamps in total?

OpenStudy (saifoo.khan):

Yes, i solved it mentally. Yes 12 stamps.

OpenStudy (anonymous):

Thank you. :) for your help.

OpenStudy (dumbcow):

you could also just graph the equation as a line...there are infinite solutions, however you need an integer solution \[y = \frac{129-16x}{7}\]

OpenStudy (lgbasallote):

there are infinite possibilities then

OpenStudy (dumbcow):

haha well mathematically yes, but practically there are finite number of positive integer solutions

OpenStudy (anonymous):

I agree. This question was abit confusing for me though.